# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 5.1. Datasets
Tables 3, 4 and 5 represent a summary of the papers that introduce datasets annotated for detecting factuality, bias and subjectivity respectively. As can be seen, the prevalent majority of the datasets are only available in English. Among the non-English for detecting factuality, subjectivity and bias, there are datasets in Urdu (Mukund and Srihari, 2010), Arabic (Mourad and Darwish, 2013) and German (Atalla et al., 2011; Aksenov et al., 2021). Furthermore, multilingual subjectivity detection was a part of shared tasks at CheckThat! Labs of CLEF 2023 (Barrón-Cedeño et al., 2023) and CLEF 2024 (Barrón-Cedeño et al., 2024), covering Arabic, German, English, Italian, and Turkish languages.
The ontology proposed by Qian et al. (2019) is the most widely accepted classification for detecting event factuality, with subsequent extension of the dataset with document-level event factuality annotations (Qian et al., 2022). More recently, Li et al. (2024c) built the first large-scale annotated dataset based on this classification, by using LLM predictions with human judgements as a final step.
Dataset | Language(s) | # Instances | Classes | Content type | |||||||||||
@@
Saurí and Pustejovsky (2009) | English | 9,488 events |
TimeBank corpus (Pustejovsky et al., 2003) | |||||||||||
@@
Li et al. (2024c) | English | 112,276 events |
Documents and events |
Dataset | Language(s) | # Instances | Classes | Content type | |||||||||||
@@
1592 news articles |
Spinde et al. (2021d) | English | 1,700 | Biased vs non-biased |
Spinde et al. (2021b) | English | 3,700 |
Liu et al. (2019a) | English | 2,990 |
Fan et al. (2019) | English | 300 news articles |
Chen et al. (2020) | English | 6964 news articles |
Aksenov et al. (2021) | German | 47362 news articles |
Dataset | Language(s) | # Instances | Classes | Content type | |||||||||
Piskorski et al. (2023) |
1592 news articles |
@@
News articles, tweets | |||||||||
Biyani et al. (2014) | English | 700 | Subjective vs Non-subjective |
@@
News articles from BBC Urdu | |||||||||
@@
Pang and Lee (2004) | English | 1,000 | Subjective vs Non-subjective | Movie reviews |
@@ 5.2. Methods and models
The methods discussed in this section can be categorized into three groups – factuality, subjectivity and bias/framing detection techniques. The methods in the first group can be further divided into two subcategories: (i) those targeting the factuality of individual event mentions and (ii) those assessing the document-level factuality of events.
@@
Subjectivity. Subjectivity detection is s well-established NLP task that emerged long before the advent of deep learning and transformer models. As a result, the vast majority of the reviewed methods rely on traditional machine learning models, such as Multinomial NaiveBayes, Convolution Kernels, Support Vector Machines, Logistic regression, Latent Dirichlet Allocation and Decision Trees over lexical, syntactic and semantic features (Biyani et al., 2014; Lin et al., 2011; Cortis and Davis, 2022; Wiegand and Klakow, 2011; Wang and Liu, 2011; Banea et al., 2010; Atalla et al., 2011; Mukund and Srihari, 2010; Mourad and Darwish, 2013; Mogadala and Varma, 2012). Rule- and lexicon-based approaches are also common for this task (Ansar and Goswami, 2021; Wiebe et al., 2004; Hammer et al., 2014; Jeronimo et al., 2020; Maks and Vossen, 2012). The only transformer-based method identified through the systematic review is that by Savinova and Del Prado (2023). The authors test the efficiency of RoBERTa model in predicting subjectivity. They found a very high correlation with the human annotators and a significantly better performance that the existing rule-based and regression methods.
Some studies formulate the task of subjectivity detection as a part of the sentiment analysis task, where the ”neutral” sentiment corresponds to the non-subjective class, and both ”positive” and ”negative” sentiments indicate subjective information (Chaturvedi et al., 2018). Barbosa and Feng (2010) explore traditional ML approached to perform the tweet sentiment prediction as a two-step approach. The authors first distinguish between subjective and non-subjective tweets, and then further perform ”positive and ”negative” tweet prediction for subjective tweets.
Finally, Lu et al. (2024) address the task of sarcasm detection, which is often seen as a specialised type of subjective information. The authors propose a novel multimodal Fact-Sentiment Incongruity Combination Network (FSICN) approach for that integrates the factual similarity and the sentiment information into the sarcasm detection task. The FSICN method incorporates a dynamic connection component that identifies the most relevant image-text pairs to detect fact incongruities between them.
Bias. The methods used for the task of bias detection can be categorized based on the type of bias concerned. At a high level, the task can be approached as a binary classification of biased vs non-biased information. The methods falling into this group typically involve traditional machine learning models and lexicon-based methods (Spinde et al., 2021d; Aggarwal et al., 2020; Hamborg et al., 2020; Lazaridou et al., 2017; Yan et al., 2019). Over the last years, transformer models and LLMs became state-of-the-art approaches towards this task (Spinde et al., 2021b; Lin et al., 2024; Wessel and Horych, 2024; Liu et al., 2019a; Fan et al., 2019; Maab et al., 2023; Krieger et al., 2022; Benson and Cruickshank, 2024).
@@
Another group of fine-grained bias detection methods concerns the idea of framing, where frames are seen as perspectives used to discuss the same topic. While some of the frames can represent political leaning, they can also be seen as discussion sub-topics. Liu et al. (2019a) analyse the dataset of news headlines on the topic of gun violence in the United States. Within this topic, they predict 9 discussion frames, such as mental health, politics, ethnicity, economical consequences, legal regulations, gun rights, public space safety and society. The authors found that BERT model significantly outperforms deep learning aprroaches, such as RNNs, LSTMs, Bi-LSTMs and Bi-GRU, for each of 9 frames.
@@ 5.3. Tools and services
Media Bias/Fact Check (MBFC) website151515https://mediabiasfactcheck.com/ is a widely used service for media factuality and bias detection. It provides systematic human-centered annotations, spanning over 8 years, regarding the factuality and bias of over 2,000 news websites. The annotators affiliated with International Fact-Checking Network (IFCN) conclude factuality in terms of a score between 0 and 10, which is in turn mapped into a 6-level scale, indicating a ”very low”, ”low”, ”mixed”, ”mostly factual”, ”high” and ”very high” rankings. To detect bias, the annotators perform an aggregation of 4 factors: biased wording, factual sourcing, story choice leaning (e.g., pro-liberal or pro-conservative), and political affiliation. Each of these categories is rated by experts on a scale of 0 to 10, from least to most biased. The final score is then calculated as a sum of the four respective scores and is mapped to the categories indicating a range of categories between extreme and moderate left and right bias and unbiased/centered reporting. The original goal of the service was to educate the public on media factuality bias and deceptive news practices. In addition, the API makes the tool highly useful for NLP classification tasks (Baly et al., 2018, 2020a; Barrón-Cedeño et al., 2023; Aires et al., 2020b; Sales et al., 2021; Kangur et al., 2024).
@@
AllSides Media Bias Ratings161616https://www.allsides.com/ is a service similar to MBFC in terms of covering the political leaning bias based on a scale from right to left. The source coverage is also comparable to MBFC, with over 2,400 websited analysed. However, unlike MBFC, the judgement is not limited to a closed group of experts, but also includes ordinary people across the political spectrum trained to perform the analysis. This provides useful insights into how people with certain political leaning biases rate media news without knowing the source. Additionally, there is an option of a community feedback to agree or disagree with the provided rating. The labelled data from AllSides has been used in a number of classification tasks for political bias detection (Baly et al., 2020a; Spinde et al., 2021a; Li and Goldwasser, 2021; Spinde et al., 2021a, d, c).
@@
Ad Fonted Media171717https://adfontesmedia.com/ provides an interactive media bias chart where media sources are annotated based on two different dimensions, reliability and political bias. The bias dimension consists of most extreme left/right, hyperpartisan left/right, strong left/right, skews left/right and middle or balanced bias. The reliability dimension consists of numerical scores that are mapped to 8 categories: fabricated news, misleading information, selective or incomplete story/propaganda, opinionated of highly varied reliability news, simple fact reporting, and high effort original fact reporting. This resource is also commonly used by NLP approaches to analyse media bias with various degrees of granularity (Chen et al., 2020; Jaradat et al., 2024).
@@
Finally, The University of Sheffield as a part of their GATE Cloud infrastructure (Cunningham et al., 2002) provides public tools for multilingual news genre and framing detection based on its submission (Wu et al., 2023) to 2023 SemEval Shared Task 3 (Piskorski et al., 2023). The models used in these tools achieved competitive results in the shared task, placing in top 3 for most of these languages. The genre detection tool181818https://cloud.gate.ac.uk/shopfront/displayItem/news-genre-classifier tackles the task of classifying the news into opinionated, objective and satire. The framing detection tool191919https://cloud.gate.ac.uk/shopfront/displayItem/news-framing-classifier allows the analysis of the 9 principal frames used to present the information.
@@ 5.4. Discussion
Lack of unified definitions and approaches. One of the main challenges of this category of credibility signals is a high diversity in how researchers perceive and address factuality, subjectivity and bias, resulting in many definitions and ontologies. This problem strengthened by the inherent subjective nature of the factuality/bias is present already in dataset annotation causing also a low inter-annotator agreement (Spinde et al., 2023). Additionally, the problem of factuality detection is often seen as a sentiment detection (Mourad and Darwish, 2013) or a hate speech detection task (Lin et al., 2022). The problem of bias detection is sometimes seen as a specific case of fake news detection, where the biased presentation of facts is automatically seen as a fake information (Lin et al., 2022). Finally, factuality, subjectivity and bias is a cumulative notion that is highly nuanced depending on how balanced the opinionated information is and how it is quoted and presented in text (Piskorski et al., 2023).
Data scarcity. The biggest challenge in multilingual subjectivity and bias detection is data scarcity. Many English studies are US-centric and only few datasets are available in other languages. For this reason, the multilingual datasets are not well-representative for benchmarking models in global analyses and multicultural settings, since bias detection requires cultural context and background knowledge about a country’s political spectrum. For instance, the concepts of right- and left-leaning media can be different across European media and Middle East media. It is important for constructing more comprehensive datasets to benchmark multilingual models.
@@ 6. Persuasion techniques and logical fallacies
Over the past decade, propagandistic efforts have been widely used on social media platforms to shape public opinion and drive engagement on a massive scale. To address this issue, several studies have proposed computational methods aimed at automatically detecting such content (Da San Martino et al., 2021). Within the scope of propagandistic content, persuasion techniques and logical fallacies aim to deliberately influence others’ opinions using rhetorical and psychological mechanisms (Miller, 1939). Recently, the task of automatic detection of persuasion techniques has gained increased attention from the NLP research community, with several resources introduced over the years.

### v2

@@ 5.1. Datasets
| Dataset | Language(s) | # Instances | Content type | Classes | |||||||||||
@@
| Li et al. (2024c) | English | 112,276 events | Documents and events |
| Dataset | Language(s) | # Instances | Content type | Classes | |||||||||
@@
1,592 news articles |
@@
News articles, tweets |
| Biyani et al. (2014) | English | 700 |
Subjective vs Non-subjective | |||||||||
@@
News articles from BBC Urdu |
@@
| Pang and Lee (2004) | English | 1,000 | Movie reviews | Subjective vs Non-subjective |
| Dataset | Language(s) | # Instances | Content type | Classes | |||||||||||
| Piskorski et al. (2023) |
1,592 news articles |
| Spinde et al. (2021d) | English | 1,700 |
Biased vs non-biased | |||||||||||
| Spinde et al. (2021b) | English | 3,700 |
| Liu et al. (2019a) | English | 2,990 |
| Fan et al. (2019) | English | 300 news articles |
| Chen et al. (2020) | English | 6,964 news articles |
| Aksenov et al. (2021) | German | 47,362 news articles |
Tables 3, 4, and 5 represent a summary of the papers that introduce datasets annotated for detecting event factuality, subjectivity, and bias respectively. As can be seen, the prevalent majority of the datasets are only available in English. Besides English, there are datasets in Urdu (Mukund and Srihari, 2010), Arabic (Mourad and Darwish, 2013) and German (Atalla et al., 2011; Aksenov et al., 2021). Furthermore, multilingual subjectivity detection was a part of shared tasks at CheckThat! Labs of CLEF 2023 (Barrón-Cedeño et al., 2023) and CLEF 2024 (Barrón-Cedeño et al., 2024), covering Arabic, German, English, Italian, and Turkish languages.
The ontology proposed by Qian et al. (2019) is the most widely accepted classification for detecting event factuality, with subsequent extension of the dataset with document-level event factuality annotations (Qian et al., 2022). More recently, Li et al. (2024c) built the first large-scale annotated dataset based on this classification, by using LLM predictions with human judgments as a final step.
@@ 5.2. Methods and models
The methods discussed in this section can be categorized into three groups – event factuality, subjectivity and bias/framing detection techniques. The methods in the first group can be further divided into two subcategories: (i) those targeting the factuality of individual event mentions and (ii) those assessing the document-level factuality of events.
@@
Subjectivity. Subjectivity detection is a well-established NLP task that emerged long before the advent of deep learning and transformer models. As a result, the vast majority of the reviewed methods rely on traditional machine learning models, such as Multinomial Naïve Bayes, Convolution Kernels, Support Vector Machines, Logistic Regression, Latent Dirichlet Allocation and Decision Trees over lexical, syntactic and semantic features (Biyani et al., 2014; Lin et al., 2011; Cortis and Davis, 2022; Wiegand and Klakow, 2011; Wang and Liu, 2011; Banea et al., 2010; Atalla et al., 2011; Mukund and Srihari, 2010; Mourad and Darwish, 2013; Mogadala and Varma, 2012). Rule- and lexicon-based approaches are also common for this task (Ansar and Goswami, 2021; Wiebe et al., 2004; Hammer et al., 2014; Jeronimo et al., 2020; Maks and Vossen, 2012). The only transformer-based method identified through the systematic review is that by Savinova and Del Prado (2023). The authors test the efficiency of RoBERTa model in predicting subjectivity. They found a very high correlation with the human annotators and a significantly better performance than the existing rule-based and regression methods.
Some studies formulate the task of subjectivity detection as a part of the sentiment analysis task, where the “neutral” sentiment corresponds to the non-subjective class, and both “positive” and “negative” sentiments indicate subjective information (Chaturvedi et al., 2018). Barbosa and Feng (2010) explore traditional ML approached to perform the tweet sentiment prediction as a two-step approach. The authors first distinguish between subjective and non-subjective tweets, and then further perform “positive” and “negative” tweet prediction for subjective tweets.
Finally, Lu et al. (2024) address the task of sarcasm detection, which is often seen as a specialised type of subjective information. The authors propose a novel multimodal Fact-Sentiment Incongruity Combination Network (FSICN) approach that integrates the factual similarity and the sentiment information into the sarcasm detection task. The FSICN method incorporates a dynamic connection component that identifies the most relevant image-text pairs to detect fact incongruities between them.
Bias. The methods used for the task of bias detection can be categorized based on the type of bias concerned. At a high level, the task can be approached as a binary classification of biased vs non-biased information. The methods falling into this group typically involve traditional machine learning models and lexicon-based methods (Spinde et al., 2021d; Aggarwal et al., 2020; Hamborg et al., 2020; Lazaridou et al., 2017; Yan et al., 2019). Over the last years, transformer models and LLMs became state-of-the-art approaches towards this task (Spinde et al., 2021b; Lin et al., 2024b; Wessel and Horych, 2024; Liu et al., 2019a; Fan et al., 2019; Maab et al., 2023; Krieger et al., 2022; Benson and Cruickshank, 2024).
@@
Another group of fine-grained bias detection methods concerns the idea of framing, where frames are seen as perspectives used to discuss the same topic. While some of the frames can represent political leaning, they can also be seen as discussion sub-topics. Liu et al. (2019a) analyse the dataset of news headlines on the topic of gun violence in the United States. Within this topic, they predict 9 discussion frames, such as mental health, politics, ethnicity, economical consequences, legal regulations, gun rights, public space safety and society. The authors found that BERT model significantly outperforms deep learning approaches, such as RNNs, LSTMs, Bi-LSTMs and Bi-GRU, for each of 9 frames.
@@ 5.3. Tools and services
Media Bias/Fact Check (MBFC) website141414https://mediabiasfactcheck.com/ is a widely used service for media factuality (combining factual accuracy with other factors) and bias detection. It provides systematic human-centered annotations, spanning over 8 years and covering over 2,000 news websites. The annotators affiliated with International Fact-Checking Network (IFCN) assess factuality based on 4 factors: failed fact checks, sourcing, transparency, and one-sidedness/omission, each rated on a scale of 0 to 10. The final factuality score (also on a scale of 0 to 10) is a weighted average of these factors. This score is further mapped into a 6-level scale, indicating a “very low”, “low”, “mixed”, “mostly factual”, “high” and “very high” rankings. To detect bias, the annotators perform an aggregation of 4 factors: biased wording, factual sourcing, story choice leaning (e.g., pro-liberal or pro-conservative), and political affiliation. Each of these categories is rated by experts on a scale of 0 to 10, from least to most biased. The final score is then calculated as a sum of the four respective scores and is mapped to the categories indicating a range of categories between extreme and moderate left and right bias and unbiased/centered reporting. The original goal of the service was to educate the public on media factuality and bias and on deceptive news practices. In addition, the API makes the tool highly useful for NLP classification tasks (Baly et al., 2018, 2020a; Barrón-Cedeño et al., 2023; Aires et al., 2020b; Sales et al., 2021; Kangur et al., 2024).
@@
AllSides Media Bias Ratings151515https://www.allsides.com/ is a service similar to MBFC in terms of covering the political leaning bias based on a scale from right to left. The source coverage is also comparable to MBFC, with over 2,400 websited analysed. However, unlike MBFC, the judgment is not limited to a closed group of experts, but also includes ordinary people across the political spectrum trained to perform the analysis. This provides useful insights into how people with certain political leaning biases rate media news without knowing the source. Additionally, there is an option of a community feedback to agree or disagree with the provided rating. The labelled data from AllSides has been used in a number of classification tasks for political bias detection (Baly et al., 2020a; Spinde et al., 2021a; Li and Goldwasser, 2021; Spinde et al., 2021a, d, c).
@@
Ad Fonted Media161616https://adfontesmedia.com/ provides an interactive media bias chart where media sources are annotated based on two different dimensions, reliability and political bias. The bias dimension consists of most extreme left/right, hyperpartisan left/right, strong left/right, skews left/right and middle or balanced bias. The reliability dimension consists of numerical scores that are mapped to 8 categories: fabricated news, misleading information, selective or incomplete story/propaganda, opinionated of highly varied reliability news, simple fact reporting, and high effort original fact reporting. This resource is also commonly used by NLP approaches to analyse media bias with various degrees of granularity (Chen et al., 2020; Jaradat et al., 2024).
@@
Finally, The University of Sheffield as a part of their GATE Cloud infrastructure (Cunningham et al., 2002) provides public tools for multilingual news genre and framing detection based on its submission (Wu et al., 2023) to SemEval 2023 Shared Task 3 (Piskorski et al., 2023). The models used in these tools achieved competitive results in the shared task, placing in top 3 for most of these languages. The genre detection tool171717https://cloud.gate.ac.uk/shopfront/displayItem/news-genre-classifier tackles the task of classifying the news into opinionated, objective and satire. The framing detection tool181818https://cloud.gate.ac.uk/shopfront/displayItem/news-framing-classifier allows the analysis of the 9 principal frames used to present the information.
@@ 5.4. Discussion
Lack of unified definitions and approaches. One of the main challenges of this category of credibility signals is a high diversity in how researchers perceive and address factuality, subjectivity and bias, resulting in many definitions and ontologies. This problem strengthened by the inherent subjective nature of the subjectivity/bias is present already in dataset annotation causing also a low inter-annotator agreement (Spinde et al., 2023). Additionally, the problem of subjectivity detection is often seen as a sentiment detection (Mourad and Darwish, 2013) or a hate speech detection task (Lin et al., 2022). The problem of bias detection is sometimes seen as a specific case of fake news detection, where the biased presentation of facts is automatically seen as false information (Lin et al., 2022). The problem of factuality detection is sometimes seen as the degree of certainty of a speaker regarding the occurrence of events, e.g., (Li et al., 2020b; Cao et al., 2021; Rovera et al., 2025) and sometimes as factual accuracy (veracity) (Nakov et al., 2024) or a mixture of veracity and other credibility signals (as is the case of the MBFC service). Finally, factuality, subjectivity and bias is a cumulative notion that is highly nuanced depending on how balanced the opinionated information is and how it is quoted and presented in text (Piskorski et al., 2023).
Data scarcity. The biggest challenge in multilingual subjectivity and bias detection is data scarcity. Many English studies are US-centric and only few datasets are available in other languages. For this reason, the multilingual datasets are not well-representative for benchmarking models in global analyses and multicultural settings, since bias detection requires cultural context and background knowledge about a country’s political spectrum. For instance, the concepts of right- and left-leaning media can be different across European or Middle East media.
@@
Adoption of LLMs. The use of large language models for detecting factuality, subjectivity, and bias is still in its early stages, largely due to two key limitations: the implicit biases embedded within LLMs themselves, and their tendency to hallucinate or generate factually incorrect information.
Recently, the task of subjectivity detection has been the focus of two large-scale studies examining it from the perspective of large language models. Suwaileh et al. (2025) conducted a comprehensive benchmark of both mono- and multilingual models for subjectivity detection in Arabic sentences. Their findings show that LLMs outperform smaller multilingual fine-tuned models even in zero-shot settings. Performance further improves when LLMs are given few-shot examples. Similarly, Shokri et al. (2024) evaluated transformer-based and large language models on subjectivity detection across three diverse datasets and domains. While fine-tuned models generally outperformed zero-shot LLMs, the latter still achieved performance close to that of the smaller fine-tuned models. The authors also explored various prompting strategies and chain-of-thought techniques, observing that in-context learning is not consistently robust, as its effectiveness is highly sensitive to the specific examples used in the prompt.
The task of bias detection has primarily been approached through the lens of framing analysis using LLMs. Pastorino et al. (2024) conducted the first benchmarking study of LLMs on framing detection, evaluating their performance in zero-shot, few-shot, and explainable prompting settings. The few-shot setting was further analysed using both in-domain and cross-domain examples. The authors found that prompting LLMs to generate explanations alongside predictions led to more robust and reliable results. While some LLMs performed particularly well in few-shot settings, the models often perceive emotional style as framing, which poses a significant challenge for reliably applying LLMs to this task. An additional challenge in applying LLMs to political bias detection is that LLMs themselves often exhibit inherent biases, which can shift in response to subtle changes in phrasing or context (Lunardi et al., 2024; Bang et al., 2024). Strategies such as prompt engineering and fine-tuning for debiasing (Lin et al., 2024a), as well as increasing model transparency regarding bias-related behaviour (Mohanty, 2025), may enhance the reliability of LLMs for future applications in this area.
Mujahid et al. (2025) explored the application of LLMs to assess political bias and factuality, where factuality is interpreted as factual accuracy (i.e., veracity) rather than as a contextual credibility signal. The authors propose two prompt-based approaches: a handcrafted method and a systematic profiling method. In the handcrafted approach, LLMs are presented with a series of questions designed to elicit judgments about a domain’s stance on specific political figures or topics, as well as assessments of the domain’s overall trustworthiness. The systematic approach involves using LLMs to profile editorial political bias across 16 policy areas. In both cases, the models were also prompted to provide justifications for their judgments. The outputs generated by the LLMs were then used as features to train traditional machine learning models and transformer-based models for predicting degrees of factuality (i.e., factual accuracy) and political bias. The study also compared these prompt-based methods to a zero-shot baseline, in which the LLM receives only the domain name, or a domain name plus a summary of five articles from the domain. Results show that the supervised prompt-based approaches significantly outperform the zero-shot baselines in both bias and factuality (i.e., factual accuracy) prediction tasks.
As mentioned above, factuality is also one of ubiquitous problems when evaluating the quality of LLMs. Research in this direction is related to how grounded LLMs’ replies are to real-world facts rather than applying LLMs to estimate factuality of information. Augenstein et al. (2024) summarised major challenges of large language models in this directions, which include unreliable citations, hallucinations when grounding information, domain-specific unreliability, incoherent responses, factual inaccuracies and difficulties performing deductive and inductive inferences. Additionally, factual accuracy of LLMs is affected by the knowledge gaps, which, depending on the information cut-off dates, can result in factually incorrect replies. These challenges become particularly dangerous when accompanied with confident tone and highly persuasive first-person replies. To address these challenges, the research community must consider a set of techniques, such as alignment of LLMs with certain values and safeguards, retrieval-based generation (RAG) to mitigate ungrounded replies, timely knowledge update and mitigation of hallucination. Finally, better evaluation metrics addressing the factuality (i.e., factual accuracy) of LLMs’ outputs need to be introduced to ensure the quality of the generated replies. Certain benchmarks have already been introduced to perform the factuality assessment of LLMs and facilitate reliable and accurate replies (Iqbal et al., 2024; Wang et al., 2024c).
@@ 6. Persuasion techniques and logical fallacies
Over the past decade, propagandistic efforts have been widely used on social media platforms to shape public opinion and drive engagement on a massive scale. To address this issue, several computational methods aimed at automatic detection of such a content have been proposed (Da San Martino et al., 2021). Within the scope of propagandistic content, persuasion techniques and logical fallacies aim to deliberately influence others’ opinions using rhetorical and psychological mechanisms (Miller, 1939). Recently, the task of automatic detection of persuasion techniques has gained increased attention from the NLP research community, with several resources introduced over the last years.